Reports 1-1 of 1 Clear search Modify search
AdV-DET (Optical benches layout)
bertolini, flaminio, gouaty, tacca, boldrini, pinto, ruggi, magazzu, gherardini - 23:37 Monday 31 May 2021 (51979) Print this report
SNEB upgrade for 40 W of input power

We implemented on SWEB the modifications in order to handle 40 W of input power:

  • with the bench suspended in air, we added some masses and both the angular and the position controls have been properly closed, after having recentered the suspension;
  • at the beginning we were not able to find any flash on B7 camera and photodiode, even if the mirrors have been realigned in the last good working point;
  • Paolo reminded that, venting the minitower and the corresponding pipe, induces a tilt on the long tower;
  • tilting about 25 urad in TX the end mirror allowed us to recover flashes and the lock of the North arm;
  • the lock of the cavity was working properly also in reflection on B4;
  • at 14 UTC some minutes of clean data have been acquired to have reference positions on B7 camera;
  • we blocked the bench to install/replace/remove the following components:
    • SNEB_M1: mirror replacement -> V213 -> W218 (R = 0.20, arrow: horizontal, on the right looking att he mirror, pointing on West);
    • SNEB_M2: mirror replacement -> V212 -> AdV206 (R = 0.85, arrow: horizontal, on the right looking att he mirror, pointing on East);
    • SNEB_M3: new mirror installation -> AdV238 (R = 0.97, arrow: horizontal, on the right looking att he mirror, pointing on West);
    • SNEB_Mmot4:  mirror replacement V214 -> AdV205 (R = 0.85, arrow: horizontal, on the right looking att he mirror, pointing on East);
    • a new beam dump on the reflection of SNEB_M3 has been installed;
    • B7_PD1 and B7_PD2 have been swapped (now Airbox 29 in reflection and airbox 30 in transmission);
    • new absorbing glasses have been added on the two B7 photodiodes;
    • one large counterweight has been removed;
    • the small absorbing glass used to dump the reflection of B7_QD2 has been removed;
    • the large absorbing glass used to dump the reflection of B7_QD2 has been replaced between QD2 and M3;
    • two counterweights have been added.
  • we unblocked the bench and we balanced it;
  • both angular and position loops have been properly closed;
  • we updated the configuration file and the offsets of the two photodiodes;
  • we roughly aligned the beam on B7_PD1 and B7_PD2.

In the following table the mass of installed and removed components is listed:


Added mass (g) Removed mass (g)
Mirror mount SNEB_M3 + M4 screw 235.6  
Mirror SNEB_M3 R = 0.2 47.3  
New beam dump 310.3  
Counter weight + screw   934.6
Two clamps for beam dump + screws 120  
One QPD beam dump + screws   176.9
One counter weight + screw 353.6  
One counter weight + screw 88.4  
Two PD beam dumps 22.5  

Missing activities:

  • centering of the B7 camera;
  • alignment of the beam on B7_QD2;
  • alignment of the beam reflected by B7_QD2 on the proper beam dump;
  • alignment of the beam reflected by SNEB_M3 on the new beam dump.

In hte attached picture the new mirror SNEB_M3 and the new beam dump are shown.

We profited of the intervention on SNEB to check also possible interferences on the bottom part of the mSAS and on the bench itself (possible sources of some of the troubles encountered in the last weeks):

  • the LVDT connector of the filter 1 is free to move and is far from the shield (rubbing almost impossible);
  • one large white cable on the bottom of the bench was rubbing against the exernal frame (it could be the source of rubbing solved slightly moving the bench along z).

Reminder for future interventions on SNEB and SWEB with the bech suspended in air: add a tilt of about 25 urad in TX to recover the proper alignment of the cavities.

Images attached to this report
Comments to this report:
tacca, berni - 17:15 Saturday 05 June 2021 (52050) Print this report

The implementation of the new SNEB configuration has been completed:

  • the North cavity has been locked on B7 in air;
  • the beam has been almost centered on the camera and the photodiodes moving both the bench and the picomotors;
  • the North cavity has been locked on B4;
  • the beam has been roughly centered on the quadrant B7_QD2: the beam in not clippend on M3, light is arriving on the sensors, it was difficult to understand the exact meaning of the B7_QD2 signals (to be further checked);
  • the beam reflected by M3 has been centered on the new beam dump installed last Monday;
  • the beam reflected by the quadrant has been sent on the corresponding beam dump;
  • the balancing of the bench has been checked;
  • the microtower has been closed and pumped up to a decent level of vacuum (slighlty lower than 1 mbar) -> some more pumping should be done during the next maintenance or whenever is possible;
  • the mSAS has been recentered in vacuum;
  • angular and position loops have been closed in vacuum;
  • the North arm has been re-aligned and re-locked in vacuum wihtout troubles.
Search Help